Lane invested $10,000 part at 15% and part at 14% If the total interest at the end of the year is 1,430, how much did she invest at 15%?
Answer by ankor@dixie-net.com(22740)

Lane invested $10,000, part at 15%, and part at 14%. If the total interest at the end of the year is 1,430, how much did she invest at 15%?
:
We know the two investment total is $10,000, therefore:
:
Let x = amt of the 15% investment
Then [10000-x] = amt of 14% investment
:
A simple interest equation:
:
.15x + .14(10000-x) = 1430
.15x + 1400 - .14x = 1430
.01x = 1430-1400
x = 30/.01
x = 3000 is the 15% amt
:
Then the 14% amt would have to be $7000, obviously
:
Check: .15(3000) + .14(7000) =
450 + 980 = 1430